Five times the width of a rectangle is 10 more than its length. If the perimeter of the rectangle is 112 yd, find the dimensions of the rectangle
Answer by checkley77(12844)

5W=L+10 OR L=5W-10
P=2L+2W
112=2(5W-10)+2W
112=10W-20+2W
112=12W-20
12W=112+20
12W=132
W=132/12
W=11 ANS. FOR THE WIDTH.
L=5*11-10
L=55-10
L=45
PROOF:
112=2*45+2*11
112=90+22
112=112